Two liquids A and B at temperatures 75 celcius and 15 celcius their masses are in ratio 2:3 and their specific heat in ratio 3:4 .What is their resulting temperature ?

Grade:12

1 Answers

Adarsh
6 years ago
Let the final temperature of mixture be T..
A/Q,    
       ma.sa.(75-T)=mb.sb.(T-15)                          
      ma/mb.sa/sb. (75-T)=(T-15)
    on putting the values we get T = 350C.
   Thanks and Regards 
            Adarsh